OpenStudy (anonymous):

The line passes thru the points (2,-5) and (8,-5)

hartnn (hartnn):

good to know :) but what do u need with that ? find equation of line ? or just slope ?

OpenStudy (anonymous):

Says to write an equation of a line given the following information.

hartnn (hartnn):

so, lets start with the slope, do u know how to find slope given 2 points ?

OpenStudy (anonymous):

i had written it \[m=\frac{ -2-(-5) }{ 8-2}\]

hartnn (hartnn):

naah, you see that y co-ordinates of both points are same, right ? so numerator of slope = y2-y1 = -5 -(-5) = -5 +5 = 0 so the entire slope = 0 got this ?

hartnn (hartnn):

then the equation would be just y=c , (because in y=mx+c, m=0) and the y co-ordinate is just -5 so the equation is just y=-5 :)

OpenStudy (anonymous):

-.-; I wrote it incorrectly on my homework when I was trying to solve it but wrote it correctly here, oy. Thank you

hartnn (hartnn):

welcome ^_^